# Time-resolved magneto-optical Kerr effect in the altermagnet candidate MnTe

### v1

@@ Abstract
Figure 4: Temperature dependence of the 3.6 THz and 4.2 THz optical phonons. (a) $\Delta\theta_{K}(t)$ traces from 10 K to 350 K and (b) corresponding FFTs. At higher temperatures, the phonons broaden in linewidth, redshift, and have shortened lifetime, in agreement with previous Raman studies. (c) Fit time scales $\tau_{1}$ and $\tau_{2}$ after fitting the background to a two-temperature model described in Eqn. 4. (d) Amplitudes and (e) frequencies of the two phonons, obtained by fitting to two damped sinusoids after background subtraction.
@@
Eqn. 5 approximates a two-temperature model in which $\Delta T_{el}(t)$ is the electron temperature, $\Delta T_{l}(t)$ is the lattice temperature, $\tau_{ee}$ is the electron-electron equilibration time, $\tau_{el}$ is the electron-phonon relaxation time, $\tau_{0}\sim$ 100 fs is the pump pulse width, and $A$ and $B$ are fitting amplitudes relating temperature changes to optical response. Since $\Delta\theta_{K}(t)$ is not identical to reflectivity, we fit to:
@@
$\begin{split}\Delta\theta_{K}(t)&=\Delta\theta_{K,1}(t)+\Delta\theta_{K,2}(t)\\ &=C_{1}\mathrm{erf}\left(\frac{t}{\tau_{1}}\right)e^{-t/\tau_{2}}+C_{2}\mathrm{erf}\left(\tau/\tau_{0}\right)\left(1-e^{-\frac{\tau+\tau_{0}}{\tau_{2}}}\right)\end{split}$ (6)
@@
without explicitly identifying $\tau_{1}$ and $\tau_{2}$ with $\tau_{ee}$ and $\tau_{el}$. Fitting results from the electron and lattice time scales are shown in Fig. 4(c). $\tau_{1}$ increases with increasing $T$, while $\tau_{2}$ decreases. Although the background is well-fit by this model (see the supplemental information for details), $\tau_{1}$ is too long to be straightforwardly associated with $\tau_{ee}$, which is typically $<$ 1 ps. However, $\tau_{2}$ is on the same order as typical $\tau_{el}$, and the increase of $\tau_{2}$ with $T$ may correspond to the known increase of $\tau_{el}$ with increasing electron temperature Mansart et al. (2013). Unlike Ref. Bossini et al., 2021, we do not observe a sublattice demagnetization time scale that changes dramatically above $T_{N}$, which suggests that the 3.6 and 4.2 THz phonons are unrelated to the tens of GHz spin oscillations.
@@
After fitting the background, we fit the residual optical phonon oscillations to two damped sinusoids as in Eqn. 1: $\Delta\theta_{K}(t)=A_{1}e^{-\alpha_{1}t}\sin(2\pi f_{1}t+\phi_{1})+A_{2}e^{-\alpha_{2}t}\sin(2\pi f_{2}t+\phi_{2})$. The phonon fit amplitudes $A_{1},A_{2}$ and frequencies $f_{1},f_{2}$ are shown in Fig. 4(d) and (e). The frequencies monotonically decrease with increasing $T$, in agreement with previous Raman studies Zhang et al. (2020a). We find that the amplitude of the 4.2 THz phonon $A_{1}$ decreases slightly with increasing $T$, while $A_{2}$ (the 3.6 THz phonon) increases until it remains constant above 250 K.
@@
In conclusion, we used TR-MOKE to identify an acoustic phonon-coupled magnon mode and two optical phonon modes in the antiferromagnetic semiconductor MnTe. Both the magnon and the optical phonon can be suppressed or enhanced by changing pump polarization, demonstrating a non-thermal origin. Somewhat surprisingly, we observe oscillations above $T_{N}$ = 310 K, at least up to 335 K, which reflects temperature-dependent contributions of both phonons and magnons to $\Delta\theta_{K}(t)$ and will be the subject of future work. Our results provide new insight into this unique magnetic system, and will be important as MnTe continues to be studied for spintronic and altermagnetic applications.
@@ Acknowledgements.
I.G. was mainly sponsored by the Army Research Office under the grant No. W911NF-20-2-0166. The construction of the pump-probe setup was supported by the Air Force Office of Scientific Research under award no. FA9550-22-1-0410. Q.D. was supported by the NSF EPM program under grant no. DMR-2213891 and the Vagelos Institute of Energy Science and Technology graduate fellowship. This work was also supported by the U. S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), Materials Sciences and Engineering Division (growth and materials characterization).
@@
Isaiah Gray: Conceptualization (equal); Data curation (lead); Formal analysis (lead); Methodology (equal); Validation (lead); Software (supporting); Visualization (lead); Writing - original draft (lead); Writing - review and editing (equal). Qinwen Deng: Data curation (supporting); Formal analysis (supporting); Methodology (equal); Software (equal). Qi Tian: Methodology (equal); Software (equal). Michael Chilcote: Resources (equal). Matthew Brahlek: Resources (equal); Writing - review and editing (equal). Liang Wu: Conceptualization (equal); Funding Acquisition (lead); Methodology (equal); Project Administration (lead), Supervision (lead), Writing - review and editing (equal).
@@ I Magnetic field dependence of optical phonons
In Fig. S1, we show short scans of $\Delta\theta_{K}(t)$ scans at $H_{app}$ = +2.2 kG, 0 kG, and -2.2 kG out-of-plane, with corresponding FFTs. The oscillations repeat nearly point-for-point, showing no change with field. Therefore, we ascribe this signal purely to optical phonons and not phonon-induced magnons. Vertical dashed lines are drawn at $f$ = 3.65 THz and $f$ = 4.26 THz in S1(b) as a visual aid to show that the peaks do not shift.
Figure S1: (a) $\Delta\theta_{K}(t)$ time traces and (b) FFTs showing optical phonons at 295 K at $H_{app}$ = +2.2 kG, 0 kG, and -2.2 kG out-of-plane.
@@ II Long scans at low temperature
Here, we show long scans at low temperature at $H_{app}$ = 500 G, demonstrating no spin oscillations. In Fig. S2(a), we show the full data set from Fig. 4 of the main text, which restricted the time extent to 20 ps so that the optical phonon oscillations could be more easily seen. In Fig. S2(b), we present a separate set of $\Delta\theta_{K}(t)$ scans with 133 ps time extent on the same sample, also with $H_{app}$ = 500 G. Within experimental sensitivity ($\sim 1\mu$V), $\Delta\theta_{K}(t)$ is constant after electron-lattice relaxation.
@@ III Fitting TR-MOKE background at low temperatures
In this section, we detail the fitting process for the low-temperature data in Fig. 4 of the main text. We show the fitting at 10 K; other temperatures are processed in exactly the same way. Fig. S3(a) shows the raw data at 10 K, the individual contributions from the best fit $\Delta\theta_{K,1}(t)$ and $\Delta\theta_{K,2}(t)$ from Eqn. 6, and their sum $\Delta\theta_{K,net}(t)$. In Fig. S3, we subtract the best-fit $\Delta\theta_{K_{n}et}(t)$ background from the data and fit the residuals to $\Delta\theta_{K}(t)=A_{1}e^{-\alpha_{1}t}\sin\left(2\pi f_{1}t+\phi_{1}\right)+A_{2}e^{-\alpha_{2}t}\sin\left(2\pi f_{2}t+\phi_{2}\right)$.
Figure S3:
Fitting 10 K data from Fig. 4 of the main text. (a) Fitting the background to Eqn. 6. (b) Fitting the residual oscillations (with background subtracted) to two damped sinusoids.
Figure S3:
@@ IV Details of fitting magnetic field and pump polarization dependence
Here we provide details of the fitting of the spin oscillations as well as the optical phonon oscillations in Figs. 1 and 2 of the main text. In Fig. S4, we show an example of fitting the magnetic field dependence at fixed ${}^{\circ}$ pump angle in Fig. 2(a) of the main text. We fit to two damped sinusoids in Fig. S4(a) and three damped sinusoids in S4(b). In this set, a third oscillation is not well-resolved ($f_{1}=3.9\pm 3.6$ GHz) and the quality of the fit is not improved. Data at other values of $H_{app}$ in Fig. 2(a) shows similar behavior.
We then fit the pump polarization dependence in Fig. S5. We fit the spin oscillation at ${}^{\circ}$ pump angle to two and three sinusoids in S5(a) and (b), respectively, and compare to similar fits of the oscillation at ${}^{\circ}$ pump angle in (c) and (d). In the ${}^{\circ}$ data, three frequencies, fitting to 25, 43, and 55 GHz, are better than two, while in the ${}^{\circ}$ data, the fits are nearly identical and $f_{3}=58\pm 5$ GHz is not well-distinguished from $f_{1}=51.8\pm 0.2$ GHz. This suggests that there are at least three frequencies - likely one magnon and two acoustic phonons - with different pump angle dependence. Finally, we fit the optical phonon pump dependence from Fig. 1(c) of the main text to two damped sinusoids, showing an example at ${}^{\circ}$ in Fig. S5(e). We plot the resulting fit amplitudes $A_{1}$ and $A_{2}$ as a function of pump angle in S5(f), and in turn fit those amplitudes to $\sin(3\theta)$ to test for three-fold symmetry. While $A_{1}$ is inconclusive, $A_{2}$ indeed exhibits three-fold symmetry.
Figure S4:
Fitting the data at $H_{app}$ = +2.2 kG from Fig. 2(a) of the main text. At this pump angle (${}^{\circ}$), three damped sines do not provide an improved fit compared to two damped sines.
Figure S4:
Figure S5:
Fitting polarization dependence from Fig. 1 and 2 of the main text. (a,b) Fitting the spin oscillation data at ${}^{\circ}$ pump angle to two and three damped sinusoids, respectively. (c,d) Repeating the process at ${}^{\circ}$ pump angle. While three sinusoids fit better than two at ${}^{\circ}$, no noticable improvement is seen at ${}^{\circ}$. (e) Example fit of the optical phonon data at ${}^{\circ}$ pump angle to two damped sinusoids. (f) Resulting fit amplitudes as a function of angle, together with fits to three-fold symmetry ($\sin 3\theta$).
Figure S5:

### v2

@@ Abstract
Figure 4: Temperature dependence of the 3.6 THz and 4.2 THz optical phonons. (a) $\Delta\theta(t)$ traces from 10 K to 350 K and (b) corresponding FFTs. At higher temperatures, the phonons broaden in linewidth, redshift, and have shortened lifetime, in agreement with previous Raman studies. (c) Fit time scales $\tau_{1}$ and $\tau_{2}$ after fitting the background to a two-temperature model described in Eqn. 4. (d) Amplitudes and (e) frequencies of the two phonons, obtained by fitting to two damped sinusoids after background subtraction.
@@
Eqn. 6 approximates a two-temperature model in which $\Delta T_{el}(t)$ is the electron temperature, $\Delta T_{l}(t)$ is the lattice temperature, $\tau_{ee}$ is the electron-electron equilibration time, $\tau_{el}$ is the electron-phonon relaxation time, $\tau_{0}\sim$ 100 fs is the pump pulse width, and $A$ and $B$ are fitting amplitudes relating temperature changes to optical response. Since $\Delta\theta(t)$ is not identical to reflectivity, we fit to:
@@
$\begin{split}\Delta\theta(t)&=\Delta\theta_{K,1}(t)+\Delta\theta_{K,2}(t)\\ &=C_{1}\mathrm{erf}\left(\frac{t}{\tau_{1}}\right)e^{-t/\tau_{2}}+C_{2}\mathrm{erf}\left(\tau/\tau_{0}\right)\left(1-e^{-\frac{\tau+\tau_{0}}{\tau_{2}}}\right)\end{split}$ (6)
@@
without explicitly identifying $\tau_{1}$ and $\tau_{2}$ with $\tau_{ee}$ and $\tau_{el}$. Fitting results from the electron and lattice time scales are shown in Fig. 4(c). $\tau_{1}$ increases with increasing $T$, while $\tau_{2}$ decreases. Although the background is well-fit by this model (see SI for details), $\tau_{1}$ is too long to be straightforwardly associated with $\tau_{ee}$, which is typically $<$ 1 ps. However, $\tau_{2}$ is on the same order as typical $\tau_{el}$, and the increase of $\tau_{2}$ with $T$ may correspond to the known increase of $\tau_{el}$ with increasing electron temperature Mansart et al. (2013). Unlike Ref. Bossini et al., 2021, we do not observe a sublattice demagnetization time scale that changes dramatically above $T_{N}$, which suggests that the 3.6 and 4.2 THz phonons are unrelated to the spin oscillations.
@@
After fitting the background, we fit the residual optical phonon oscillations to two damped sinusoids as in Eqn. 2: $\Delta\theta(t)=A_{1}e^{-\alpha_{1}t}\sin(2\pi f_{1}t+\phi_{1})+A_{2}e^{-\alpha_{2}t}\sin(2\pi f_{2}t+\phi_{2})$. The phonon fit amplitudes $A_{1},A_{2}$ and frequencies $f_{1},f_{2}$ are shown in Fig. 4(d) and (e). The frequencies monotonically decrease with increasing $T$, in agreement with previous Raman studies Zhang et al. (2020a). We find that the amplitude of the 4.2 THz phonon $A_{1}$ decreases slightly with increasing $T$, while $A_{2}$ (the 3.6 THz phonon) increases until it remains constant above 250 K.
@@
In conclusion, we used TR-MOKE to identify a magnon mode and two optical phonon modes in the antiferromagnetic semiconductor MnTe. Both the magnon and the optical phonon can be suppressed or enhanced by changing pump polarization, demonstrating a non-thermal origin. We continued to observe field-dependent oscillations above $T_{N}$, at least up to 335 K, which may reflect the persistence of magnetic anisotropy and coupling to short-range order above $T_{N}$ and will be the subject of future work. Our results provide new insight into this unique magnetic system, and will be important as MnTe continues to be studied for spintronic and altermagnetic applications.
@@ Acknowledgements.
The authors thank Prof. Jay Kikkawa for helpful discussions. I.G. was mainly sponsored by the Army Research Office under the grant No. W911NF-20-2-0166. The construction of the pump-probe setup was supported by the Air Force Office of Scientific Research under award no. FA9550-22-1-0410. Q.D. was supported by the NSF EPM program under grant no. DMR-2213891 and the Vagelos Institute of Energy Science and Technology graduate fellowship. Q.T. was support by the US Office of Naval Research (ONR) through the grant N00014-24-1-2064. The work at ORNL was supported by the U. S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), Materials Sciences and Engineering Division (growth and materials characterization).
@@
Isaiah Gray: Conceptualization (equal); Data curation (lead); Formal analysis (lead); Methodology (equal); Validation (lead); Software (supporting); Visualization (lead); Writing - original draft (lead); Writing - review and editing (equal). Qinwen Deng: Data curation (supporting); Formal analysis (supporting); Methodology (equal); Software (equal). Qi Tian: Methodology (equal); Software (equal). Michael Chilcote: Resources (equal). J. Steven Dodge: Writing: review and editing (supporting). Matthew Brahlek: Resources (equal); Writing - review and editing (equal). Liang Wu: Conceptualization (equal); Funding Acquisition (lead); Methodology (equal); Project Administration (lead), Supervision (lead), Writing - review and editing (equal).
@@ I Magnetic field dependence of optical phonons
In Fig. S1, we show short scans of $\Delta\theta(t)$ scans at $H_{app}$ = +2.2 kG, 0 kG, and -2.2 kG out-of-plane, with corresponding FFTs. The oscillations repeat nearly point-for-point, showing no change with field. Therefore, we ascribe this signal purely to optical phonons and not phonon-induced magnons. Vertical dashed lines are drawn at $f$ = 3.65 THz and $f$ = 4.26 THz in S1(b) as a visual aid to show that the peaks do not shift.
Figure S1: (a) $\Delta\theta(t)$ time traces and (b) FFTs showing optical phonons at 295 K at $H_{app}$ = +2.2 kG, 0 kG, and -2.2 kG out-of-plane.
@@ II Long scans at low temperature
Here, we show long scans at low temperature at $H_{app}$ = 500 G, demonstrating no magnon oscillations. In Fig. S2(a), we show the full data set from Fig. 4 of the main text, which restricted the time extent to 20 ps so that the optical phonon oscillations could be more easily seen. In Fig. S2(b), we present a separate set of $\Delta\theta(t)$ scans with 133 ps time extent on the same sample, also with $H_{app}$ = 500 G. Within experimental sensitivity ($\sim 1\mu$V), $\Delta\theta(t)$ is constant after electron-lattice relaxation.
@@ III Fitting TR-MOKE background at low temperatures
In this section, we detail the fitting process for the low-temperature data in Fig. 4 of the main text. We show the fitting at 10 K; other temperatures are processed in exactly the same way. Fig. S3(a) shows the raw data at 10 K, the individual contributions from the best fit $\Delta\theta_{1}(t)$ and $\Delta\theta_{2}(t)$ from Eqn. 6, and their sum $\Delta\theta_{net}(t)$. In Fig. S3(b), we subtract the best-fit background from the data and fit the residuals to $\Delta\theta(t)=A_{1}e^{-\alpha_{1}t}\sin\left(2\pi f_{1}t+\phi_{1}\right)+A_{2}e^{-\alpha_{2}t}\sin\left(2\pi f_{2}t+\phi_{2}\right)$.
Figure S3: Fitting 10 K data from Fig. 4 of the main text. (a) Fitting the background to Eqn. 6. (b) Fitting the residual oscillations (with background subtracted) to two damped sinusoids.
@@ IV Details of fitting magnetic field and pump polarization dependence
Here we provide details of the fitting of the spin oscillations as well as the optical phonon oscillations in Figs. 1 and 2 of the main text. In Fig. S4, we show an example of fitting the magnetic field dependence at fixed 225 pump angle in Fig. 2(a) of the main text. We fit to two damped sinusoids in Fig. S4(a) and three damped sinusoids in S4(b). In this set, a third oscillation is not well-resolved ($f_{1}=3.9\pm 3.6$ GHz) and the quality of the fit is not improved by including it. Data at other values of $H_{app}$ in Fig. 2(a) shows similar behavior.
We then fit the pump polarization dependence in Fig. S5. We fit the spin oscillation at 75 pump angle to two and three sinusoids in S5(a) and (b), respectively, and compare to similar fits of the oscillations at 195 pump angle in (c) and (d). In the 75 data, three frequencies, fitting to 25, 43, and 55 GHz, are better than two, while in the 195 data, the fits are nearly identical and $f_{3}=58\pm 5$ GHz is not well-distinguished from $f_{1}=51.8\pm 0.2$ GHz. This suggests that there are at least three frequencies - likely one magnon and two acoustic phonons - with different pump angle dependence.
Figure S4: Fitting the data at $H_{app}$ = +2.2 kG from Fig. 2(a) of the main text. At this pump angle (225), three damped sines do not provide an improved fit compared to two damped sines.
Figure S5: Fitting polarization dependence from Fig. 1 and 2 of the main text. (a,b) Fitting the spin oscillation data at 75 pump angle to two and three damped sinusoids, respectively. (c,d) Repeating the process at 195 pump angle. While three sinusoids fit better than two at 75, no noticable improvement is seen at 195.
V Fitting $\Delta\theta(t)$ above $T_{N}$
Table 1: Fit parameters for the field-dependent data in Fig. 3(b) and (c) of the main text.
$f_{1}$ (GHz) $f_{2}$ (GHz) $f_{3}$ (GHz) $A_{1}$ (mdeg) $A_{2}$ (mdeg) $A_{3}$ (mdeg)
295 K, +2.2 kG 13.7(5) 33.8(4) 59.6(1.5) 0.19(1) 0.36(5) 0.25(6)
295 K, -2.2 kG N/A 33.8(2) 64(1) N/A 0.69(5) 1.0(4)
325 K, +2.2 kG N/A 31.4(4) 58(2) N/A 0.20(3) 1.4(5)
325 K, 0 kG 15.3(3) 34.9(3) N/A 0.07(1) 0.33(4) N/A
325 K, -2.2 kG 18(2) 38.3(4) 59.1(4) 0.6(1) 0.56(5) 0.68(7)
In this section, we show the fit parameters for the field-dependent data in Fig. 3(b) and (c) of the main text. We follow the same fitting procedure as in the rest of the text: we fit the raw data to a background of a decaying exponential plus a linear function, subtract off the background, then fit the remaining oscillations to three exponentially damped sinusoids. Resulting frequencies and amplitudes for each sinusoid are shown in Table 1, with uncertainties on the last significant digit given in parentheses. As in Fig. 2 of the main text, the data at 295 K and 0 kG did not exhibit sufficient oscillation amplitude to allow a consistent fit.
We find that at both 295 K and 325 K, the $f_{3}$ mode (between 58 and 64 GHz) appears only when $H_{app}\neq$ 0, whereas the $\sim$15 GHz and 35 GHz modes are also present at 325 K and 0 kG. Note that no constraints on the frequency values were imposed in the fitting, so the presence or absence of the 55 GHz mode is not an artifact of bounding the frequencies.